Find the impedance of a series RLC circuit if the inductive reactance, capacitive
reactance and resistance are 184 Ω, 144 Ω and -j30 Ω respectively. Also calculate the
phase angle between voltage and current
1
Expert's answer
2021-10-11T08:55:26-0400

The impedance of a series RLC circuit is



"Z=\sqrt{(X_L-X_C)^2+R^2}=50\space\Omega."

The phase angle between voltage and current equals



"\phi=\text{atan}\frac{X_L-X_C}{R}=53^\circ."